# F-Term Hybrid Inflation, Metastable Cosmic Strings and Low Reheating in View of ACT

### v1

@@ Abstract:
We consider the formation of metastable cosmic strings in a left-right unified theory. The produced monopoles are diluted by a stage of F-term hybrid inflation (FHI) which is realized consistently with the SUSY breaking and a global $U(1)$ $R$ symmetry in the context of a $U(1)_{\rm R}\times U(1)_{B-L}$ extension of MSSM. The hidden sector Kähler manifold enjoys an enhanced $SU(1,1)/U(1)$ symmetry with the scalar curvature determined by the achievement of a SUSY-breaking de Sitter vacuum without ugly tuning. FHI turns out to be compatible with data – including the recent ACT results –, provided that the magnitude of the emergent soft tadpole term is confined in the range $(0.1-70)$ TeV, and it is accompanied with the production of cosmic strings. Their dimensionless tension ${G\mu_{\rm cs}}\simeq(1-11)\cdot 10^{-8}$ interprets the present observations from PTA experiments on the stochastic background of gravitational waves. The $\mu$ parameter of MSSM arises by appropriately adapting the Giudice-Masiero mechanism and facilitates the out-of-equilibrium decay of the $R$ saxion at a reheat temperature lower than about $34$ GeV. The SUSY mass scale turns out to lie in the PeV region.
@@ 1 Introduction
In this talk we attempt to bring together the recent data [1, 2] on gravitational waves (GWs), inflation [3, 4] and Supersymmetry (SUSY) breaking [5]. The junction point is the metastable cosmic strings (CSs) [6] whose decay interprets the recent data on GWs. Their generation requires a specific GUT construction – see Sec. 1.1 – in conjunction with a suitable inflationary stage – see Sec. 1.2.
@@ 1.1 PTA Data and Metastable CSs
The discovery of a background of GWs around the nanohertz frequencies announced from several Pulsar Timing Array (PTA) experiments [1] – most notably the NANOGrav 15-years results (NG15) – provides a novel tool in exploring the structure of early universe. The observations can be explained by gravitational radiation emitted by topologically unstable superheavy CSs which may be formed during the spontaneous symmetry breaking (SSB) chains of Grand Unified Theories (GUTs) [7] down to the Standard Model (SM) gauge group $\mathbb{G}_{\rm SM}$. In particular, the observations can be interpreted if the CSs are metastable. This type of CSs arise from a SSB of the form
@@
and so ${\rm adj}(\mathbb{G})$ may be identified by $({\bf 1,1,3},0)$. Since the production of MMs is cosmologically catastrophic, these are to be inflated away. However, they can appear on CSs via quantum pair creation if the formatted CSs are metastable.
@@ 1.2 CSs, Inflation and ACT Data
$\mathbb{G}_{\rm LR}\times U(1)_{R}\times U(1)_{B}\times{\rm SUGRA}{}{}$ $\underset{\rm MMs}{\xrightarrow{\hskip 21.33955pt\langle{({\bf 1,1,3},0)}\rangle=v_{\rm R}\hskip 21.33955pt}}~{}~{}\mathbb{G}_{\rm L1R}\times U(1)_{R}\times U(1)_{B}\times{\rm SUGRA}$ (5)
$-\left(\mbox{\rm\footnotesize FHI}\right)\underset{\langle{Z}\rangle,~{}\mbox{\rm\footnotesize CSs}}{{\xrightarrow{\hskip 11.38092pt\langle{\bar{\Phi}}\rangle=\langle{\Phi}\rangle=M\hskip 11.38092pt}}}~{}~{}\mathbb{G}_{\rm SM}\times\mathbb{Z}_{2}^{R}\times U(1)_{B}\times~{}~{}\leavevmode\hbox{{\Large$\diagup$}{\kern-18.0pt\normalsize SUSY}}$
$~{}\underset{\rm RCs}{{\xrightarrow{\hskip 28.16838pt\langle{{H_{d}}}\rangle,\langle{{H_{u}}}\rangle\hskip 28.16838pt}}}~{}~{}SU(3)_{\rm C}\times U(1)_{\rm EM}\times U(1)_{B},$
where we include also the final transition to the present vacuum which occurs via the well-known radiative electroweak SSB – EM stands for ElectroMagnetism. Contrary to Ref. [21, 20] we do not specify the dynamics of the SSB of $\mathbb{G}_{\rm LR}$ but we incorporate the mechanism of SUSY breaking as shown in the rightmost part of the second line. Note that soft SUSY breaking effects explicitly break $U(1)_{R}$ to the $\mathbb{Z}_{2}^{R}$ which remains unbroken and, if combined with the fermion parity, yields the well-known $R$-parity of MSSM.
@@ 2.1 Particle Content
As already mentioned, we focus on the second line of Eq. (5) and therefore we consider a model invariant under the gauge group $\mathbb{G}_{\rm L1R}$ in Eq. (4) which incorporates FHI. In addition to the local symmetry, the model possesses the baryon and lepton number symmetries and an $R$ symmetry $U(1)_{R}$. The representations under $\mathbb{G}_{\rm L1R}$ and the charges under the global symmetries of the various matter and Higgs superfields of the model are presented in Table 1. Namely, the $i$th generation $SU(2)_{\rm L}$ doublet left-handed quark and lepton superfields are denoted by $q_{i}$ and $l_{i}$ respectively, whereas the $SU(2)_{\rm L}$ singlet antiquark [antilepton] superfields by $u^{c}_{i}$ and ${d_{i}}^{c}$ [$e^{c}_{i}$ and $\nu^{c}_{i}$] respectively. The electroweak Higgs superfields which couple to the up [down] quark superfields are denoted by ${H_{u}}$ [${H_{d}}$]. Besides the MSSM particle content, the model is augmented by seven superfields: a gauge singlet $S$, three $\nu^{c}_{i}$’s, a pair of Higgs superfields $\Phi$ and $\bar{\Phi}$ which break $U(1)_{\rm R}\times U(1)_{B-L}$ and the goldstino superfield $Z$. Note that the SM hypercharge $Q_{Y}$ is identified as the linear combination $Q_{Y}=Q_{\rm R}+Q_{(B-L)}/2$ where $Q_{\rm R}$ and $Q_{(B-L)}$ is the $U(1)_{\rm R}$ and $B-L$ charge respectively.
@@ 2.3 Kähler Potential
Our favored $\nu$ range finally is $13/4<\nu<$ 1. Since $N<0$, $K_{\rm H}$ parameterizes the $SU(1,1)/U(1)$ hyperbolic Kähler manifold for $k\sim 0$.
@@ 3 Inflation Analysis
are the constant potential energy density and correspoding Hubble parameter which drive FHI – the subscript 00 means that this is the tree level value. In the present context, though, prominent role in the investigation plays the F-term SUGRA potential $V_{\rm F}$ which is minimized under the conditions displayed in Sec. 3.1 and then in Sec. 3.2 we give the final form of the inflationary potential. Lastly, we present our results in Sec. 3.4 imposing a number of constraints listed in Sec. 3.3.
@@ 3.3 Observational Requirements
The number of e-foldings elapsed from horizon exit of $k_{\star}=0.05/{\rm Mpc}$ $k$ until the end of inflation at $\sigma_{\rm f}\simeq\sigma_{\rm c}$ have to be enough to resolve the problems of the Standard Big Bang, i.e., [3]:
@@
[…]
@@ 3.4 Results
The inflationary part of our model depends on the parameters – cf. Ref. [14]$\kappa,~{}M~{}~{}\mbox{and}~{}~{}{\rm a}_{S}$, where ${\rm a}_{S}$ can be derived from $m$ and $\nu$ via the rightmost relation in Eq. (18b). Enforcing Eqs. (19) and (20) we can restrict $M$ and $\sigma_{\star}$ as functions of $\kappa$ and ${\rm a}_{S}$. On the other hand, the correct $n_{\rm s}$ values in Eq. (22) are attained if FHI becomes of hilltop type [14, 15] via a careful selection of ${\rm a}_{S}$. I.e., $V_{\rm I}$ is non-monotonic and develops a maximum at $\sigma_{\rm max}$ and a minimum at $\sigma_{\rm min}\gg\sigma_{\rm max}$. For $\sigma>\sigma_{\rm min}$, $V_{\rm I}$ becomes a monotonically increasing function of $\sigma$ and so its boundedness is assured. FHI takes place for $\sigma<\sigma_{\rm max}$. The achievement of successful hilltop FHI requires the establishment of the hierarchy $\sigma_{\rm c}<\sigma_{\star}<\sigma_{\rm max}$. As a consequence, the obtained $n_{\rm s}$ values decline from the prediction $n_{\rm s}\simeq 1-1/{N_{\rm I\star}}$ of the original model [10, 16] which provides the correct value in Eq. (22), though, for the ${N_{\rm I\star}}\sim 41$.
@@
$0.84\lesssim{M/{\rm YeV}}\lesssim 2.72,~{}~{}2\lesssim{\kappa/10^{-4}}\lesssim 14~{}~{}\mbox{and}~{}~{}0.11\lesssim{{\rm a}_{S}/{\mbox{\rm TeV}}}\lesssim 70$ (23)
@@ 4.1 SUSY and $\mathbb{G}_{\rm L1R}$ Breaking
[…]
@@ 4.2 Low Reheating
Figure 2: Allowed (shaded) regions in the $\kappa-T_{\rm rh}$ (a) and $\kappa-\widetilde{m}$ (b) plane compatible with the inflationary requirements in Sec. 3.3 for $n_{\rm s}=0.974$. We take (a) $\nu=7/8$ and $\mu=\widetilde{m}$ (solid line), $\mu=\widetilde{m}/3$ (dot-dashed line) or $\mu=3\widetilde{m}$ (dashed line) whereas the BBN lower bounds on $T_{\rm rh}$ for hadronic branching ratios ${\sf B}_{\rm h}=1$ and $0.001$ are also depicted by two thin lines (b) $\widetilde{m}/3\leq\mu\leq 3\widetilde{m}$ and $13/4<\nu<$ 1 whereas the allowed contour for $\nu=7/8$ is depicted by a solid line and the region excluded by BBN for ${\sf B}_{\rm h}=0.001$ is hatched.
@@ 4.3 SUSY-Mass Scale
$0.75<\nu<1~{}~{}\mbox{and}~{}~{}1/3\lesssim\mu/\widetilde{m}\lesssim 3$ (32)
@@
$0.16\lesssim\widetilde{m}/{\mbox{\rm PeV}}\lesssim 9.2~{}~{}\mbox{with}~{}~{}T_{\rm rh}^{\max}\simeq 34$ (33)
@@ 5 GWs from Metastable CSs
[…]
@@
$1.1\lesssim M/{\mbox{\rm YeV}}\lesssim 2.72~{}~{}\mbox{and}~{}~{}3\lesssim\kappa/10^{-4}\lesssim 14$ (38)

### v2

@@ Abstract:
We consider the formation of metastable cosmic strings in a left-right unified theory. The produced monopoles are diluted by a stage of F-term hybrid inflation (FHI) which is realized consistently with the SUSY breaking and a global $U(1)$ $R$ symmetry in the context of a $U(1)_{\rm R}\times U(1)_{B-L}$ extension of MSSM. The hidden sector Kähler manifold enjoys an enhanced $SU(1,1)/U(1)$ symmetry with the scalar curvature determined by the achievement of a SUSY-breaking de Sitter vacuum without ugly tuning. FHI turns out to be compatible with data – including the recent ACT results –, provided that the magnitude of the emergent soft tadpole term is confined in the range (0.170)(0.1-70) TeV, and it is accompanied with the production of cosmic strings. Their dimensionless tension ${G\mu_{\rm cs}}\simeq(1-11)\cdot 10^{-8}$ interprets the present observations from PTA experiments on the stochastic background of gravitational waves. The $\mu$ parameter of MSSM arises by appropriately adapting the Giudice-Masiero mechanism and facilitates the out-of-equilibrium decay of the $R$ saxion at a reheat temperature lower than about $34$ GeV. The SUSY mass scale turns out to lie in the PeV region.
Published in  PoS CORFU 2024, 206 (2025)
@@ 1 Introduction
In this talk we attempt to bring together the recent data [1, 2] on gravitational waves (GWs), inflation [3, 4] and Supersymmetry (SUSY) breaking [5]. The junction point is the metastable cosmic strings (CSs) [6] whose decay interprets the recent data on GWs. Their generation requires a specific Grand Unified Theory (GUT) construction – see Sec. 1.1 – in conjunction with a suitable inflationary stage – see Sec. 1.2.
@@ 1.1 PTA Data and Metastable CSs
The discovery of a background of GWs around the nanohertz frequencies announced from several Pulsar Timing Array (PTA) experiments [1] – most notably the NANOGrav 15-years results (NG15) – provides a novel tool in exploring the structure of early universe. The observations can be explained by gravitational radiation emitted by topologically unstable superheavy CSs which may be formed during the spontaneous symmetry breaking (SSB) chains of GUTs [7] down to the Standard Model (SM) gauge group $\mathbb{G}_{\rm SM}$. In particular, the observations can be interpreted if the CSs are metastable. This type of CSs arise from a SSB of the form
@@
and so ${\rm adj}(\mathbb{G})$ may be identified by $({\bf 1,1,3},0)$. Although not cosmologically catastrophic, the production of MMs may have an impact on the GWs if they are inflated away. This is because they can appear on CSs via quantum pair creation if the formatted CSs are metastable.
@@ 1.2 CSs, Inflation and ACT Data
[…]
$-\hskip-4.26773pt\left(\mbox{\rm\footnotesize FHI}\right)\hskip-4.26773pt\underset{\langle{Z}\rangle,~\mbox{\rm\footnotesize CSs}}{{\xrightarrow{\hskip 11.38092pt\langle{\bar{\Phi}}\rangle=\langle{\Phi}\rangle=M\hskip 11.38092pt}}}~~\mathbb{G}_{\rm SM}\times U(1)_{B}\times~~\hbox{{\Large$\diagup$}{\kern-18.0pt\normalsize SUSY}}$
$~\underset{\rm RCs}{{\xrightarrow{\hskip 28.16838pt\langle{{H_{d}}}\rangle,\langle{{H_{u}}}\rangle\hskip 28.16838pt}}}~~SU(3)_{\rm C}\times U(1)_{\rm EM}\times U(1)_{B}.$
The final transition to the present vacuum occurs via the well-known radiative electroweak SSB – EM stands for ElectroMagnetism. Contrary to Ref. [21, 20] we do not specify the dynamics of the SSB of $\mathbb{G}_{\rm LR}$ but we incorporate the mechanism of SUSY breaking as shown in the rightmost part of the second line. Note that soft SUSY-breaking effects explicitly break $U(1)_{R}$ to a discrete subgroup which is spontaneously broken by the v.e.v of the sgoldstino, $Z$, too. Thanks to the explicit violation of $U(1)_{R}$ in the Kähler potential – see below – no problem with domain walls arises.
@@ 2.1 Particle Content
As already mentioned, we focus on the second line of Eq. (5) and therefore we consider a model invariant under the gauge group $\mathbb{G}_{\rm L1R}$ in Eq. (4) which incorporates FHI. In addition to the local symmetry, the model possesses the baryon and lepton number symmetries and an $R$ symmetry $U(1)_{R}$. The representations under $\mathbb{G}_{\rm L1R}$ and the charges under the global symmetries of the various matter and Higgs superfields of the model are presented in Table 1. Namely, the $i$th generation $SU(2)_{\rm L}$ doublet left-handed quark and lepton superfields are denoted by $q_{i}$ and $l_{i}$ respectively, whereas the $SU(2)_{\rm L}$ singlet antiquark [antilepton] superfields by $u^{c}_{i}$ and ${d_{i}}^{c}$ $e^{c}_{i}$ and $\nu^{c}_{i}$] respectively. The electroweak Higgs superfields which couple to the up [down] quark superfields are denoted by $H_{u}$} $H_{d}$}]. Besides the MSSM particle content, the model is augmented by seven superfields: a gauge singlet $S$, three $\nu^{c}_{i}$’s, a pair of Higgs superfields $\Phi$ and $\bar{\Phi}$ which break $U(1)_{\rm R}\times U(1)_{B-L}$ and the goldstino superfield $Z$. Note that the SM hypercharge $Q_{Y}$ is identified as the linear combination $Q_{Y}=Q_{\rm R}+Q_{(B-L)}/2$ where $Q_{\rm R}$ and $Q_{(B-L)}$ is the $U(1)_{\rm R}$ and BLB-L charge respectively.
@@ 2.3 Kähler Potential
Our favored $\nu$ range finally is $3/4<\nu<1$. Since $N<0$, $K_{\rm H}$ parameterizes the $SU(1,1)/U(1)$ hyperbolic Kähler manifold for $k\sim 0$.
@@ 3 Inflation Analysis
are the constant potential energy density and correspoding Hubble parameter which drive FHI – the subscript 0 means that this is the tree level value. In the present context, though, prominent role in the investigation plays the F-term SUGRA potential $V_{\rm F}$ which is minimized under the conditions displayed in Sec. 3.1 and then in Sec. 3.2 we give the final form of the inflationary potential. Lastly, we present our results in Sec. 3.4 imposing a number of constraints listed in Sec. 3.3.
@@ 3.3 Observational Requirements
The number of e-foldings elapsed from horizon exit of $_{\star}=0.05/{\rm Mpc} kk$ until the end of inflation at $\sigma_{\rm f}\simeq\sigma_{\rm c}$ have to be enough to resolve the problems of the Standard Big Bang, i.e., [3]:
@@
$\sqrt{A_{\rm s}}=\frac{1}{2\sqrt{3}\pi m^{3}_{\rm P}}\;\left.\frac{V_{\rm I}^{3/2}(\sigma_{\star})}{|V^{\prime}_{\rm I}(\sigma_{\star})|}\right.\simeq\>4.618\cdot 10^{-5}.$ (20)
@@ 3.4 Results
[…]
@@
$0.84\lesssim{M/{\rm YeV}}\lesssim 2.72,~~2\lesssim{\kappa/10^{-4}}\lesssim 14~~\mbox{and}~~0.11\lesssim{{\rm a}_{S}/{\mbox{\rm TeV}}}\lesssim 70.$ (23)
@@ 4.1 SUSY and $\mathbb{G}_{\rm L1R}$ Breaking
Typical values for the masses above are given in Table 2, where we see that besides $m_{\rm I}$, which is of order $1~{\mbox{\rm ZeV}}$, the other masses are of order $1~{\mbox{\rm PeV}}$.
@@ 4.2 Low Reheating
Figure 2: Allowed (shaded) regions in the $\kappa-T_{\rm rh}$ (a) and $\kappa-\widetilde{m}$ (b) plane compatible with the inflationary requirements in Sec. 3.3 for $n_{\rm s}=0.974$. We take (a) $\nu=7/8$ and $\mu=\widetilde{m}$ (solid line), $\mu=\widetilde{m}/3$ (dot-dashed line) or $\mu=3\widetilde{m}$ (dashed line) whereas the BBN lower bounds on $T_{\rm rh}$ for hadronic branching ratios ${\sf B}_{\rm h}=1$ and $0.001$ are also depicted by two thin lines (b) $\widetilde{m}/3\leq\mu\leq 3\widetilde{m}$ and $3/4<\nu<1$ whereas the allowed contour for $\nu=7/8$ is depicted by a solid line and the region excluded by BBN for ${\sf B}_{\rm h}=0.001$ is hatched.
@@ 4.3 SUSY-Mass Scale
$0.75<\nu<1~~\mbox{and}~~1/3\lesssim\mu/\widetilde{m}\lesssim 3.$ (32)
@@
$0.16\lesssim\widetilde{m}/{\mbox{\rm PeV}}\lesssim 9.2~~\mbox{with}~~T_{\rm rh}^{\max}\simeq 34.$ (33)
@@ 5 GWs from Metastable CSs
$10^{-8}\lesssim{G\mu_{\rm cs}}\lesssim 2.4\cdot 10^{-4}~~\mbox{for}~~8.2\gtrsim\sqrt{{r_{\rm ms}}}\gtrsim 7.5~~\mbox{at 95$\%$ c.l.},$ (36)
@@
$1.1\lesssim M/{\mbox{\rm YeV}}\lesssim 2.72~~\mbox{and}~~3\lesssim\kappa/10^{-4}\lesssim 14.$ (38)
@@ 6 Conclusions
Acknowledgments
I would like to thank R. Maji and Q. Shafi for useful suggestions, M. Axenides, A. Passias and K. Tamvakis for interesting questions.